jigglypuff314 (jigglypuff314):

If Rebecca gives Chloe 8 dollars, they will have the same amount. If Chloe gives Rebecca 8 dollars, Rebecca will have 5 times as much as Chloe has. How much did each have to begin with?

OpenStudy (anonymous):

R - 8 = C + 8 R+8 = 5(C-8) solve, R = 32, C = 16

jigglypuff314 (jigglypuff314):

okay, thank you! :)